OpenStudy (anonymous):

check answer please... given the function f described by f(x)=-x-4....find f(-8)

OpenStudy (anonymous):

2?

OpenStudy (e.mccormick):

Not 2. That would be f(-6)

OpenStudy (anonymous):

oh wait i multiplied

OpenStudy (rajee_sam):

y = -x-8; put x = -8 for f(-8); y = - ( -8) - 4; y = 8-4 = 4

OpenStudy (anonymous):

i did it all wrong then. wait if i use -6 it would be 4?

OpenStudy (rajee_sam):

f(-6) will be y= - (-6)-4 = 6-4 = 2

OpenStudy (e.mccormick):

The key here is watching the signs. \(f(-8)=-(-8)-4\) Beacuse the x is replaced with the \(-8\) the sign goes along. Then the \(-(-8)\) can be said to be \((-1)\times (-8)\) which is just 8.

OpenStudy (anonymous):

dont i substitute -8 for -x?

OpenStudy (e.mccormick):

Not for -x, but for x.

OpenStudy (anonymous):

ok so i would have -8=-x-4?

OpenStudy (e.mccormick):

No, that is replacing f(x) with -8. That is the wrong place to replace it. f(x) means, "The function f with respect to the variable x."

OpenStudy (anonymous):

ok im a little confused

OpenStudy (e.mccormick):

Lets take it one part at a time. First, the function name. If it is \(f(x), g(x), h(a)\) or whatever, that part just tells you the name of the function and what variable you are dealing with. This time it is the function \(f\) and the variable \(x\).

OpenStudy (e.mccormick):

That is all the left hand side of this means. With me so far?

OpenStudy (anonymous):

ok got it.

OpenStudy (e.mccormick):

The right hand side is what equation the function represents. This time it happens to be \(-x-4\). It can be one of many, many equations, but at the moment it is just that one.

OpenStudy (anonymous):

so i have to work from the -x-4

OpenStudy (e.mccormick):

Yes. What \(f(-8)\) means when you have \(f(x)=-x-4\) is this. "Look at \(-x-4\) and any time you see \(x\), replace it with \((-8)\). And I mean including the ( ) to help make it clear.

OpenStudy (anonymous):

oh hold on let me see if i have it..

OpenStudy (anonymous):

so if i replace it, it would be -8=-x-4

OpenStudy (e.mccormick):

Still replacing the wrong part. You write it out like this: \(f(-8)=-(-8)-4\)

OpenStudy (anonymous):

ohh duhhh. see i was thinking not -x cause of it being negative duhhh

OpenStudy (e.mccormick):

Hehe. Well, it is still an x. And there is even a 1 hiding there. It can be said that \(-x-4=-1x-4\)

OpenStudy (anonymous):

ok so it would be -8=-(-8)-4? do i still put the negative sign with the 8 because theres one already there?

OpenStudy (e.mccormick):

Yes, even though there already is one, you need to bring along the entire \(-8\). Now, what happens to a \(-\) when two negative numbers are multiplied?

OpenStudy (e.mccormick):

You keep premoving the f on the left. The f stays there. You can change it to y, but that is it! So iy would be \(f(-8)=-(-8)-4\) or \(y=-(-8)-4\)

OpenStudy (anonymous):

it would be a postitive

OpenStudy (anonymous):

so i would get 32

OpenStudy (e.mccormick):

Yes! So: \(f(-8)=-(-8)-4\) means: \(f(-8)=-1\times (-8)-4\) which simplifies to: \(f(-8)=8-4\) The -8 becomes positive. Do you get that part now?

OpenStudy (e.mccormick):

Umm, still did a multiplication rather than a subtraction... not 32.

OpenStudy (anonymous):

it would be a -4,, if subtraction -8 and 4

OpenStudy (e.mccormick):

The -4 means subtract 4. That - sign stays there.

OpenStudy (anonymous):

what i keep thinking of changing the subtraction sign to a plus and the 4 become -4..is that something of what your saying?

OpenStudy (e.mccormick):

It has always been a -4. \(f(x)=-x\Large-4\) \(\leftarrow\) there it is, a -4.

OpenStudy (anonymous):

ok i was getting off on something else

OpenStudy (e.mccormick):

No problem. You are obviously just starting functions. It is a new topuc. New topics tend to be confusing.

OpenStudy (anonymous):

yea i just started them. the function machines i can do a little better with, but this problem i just dont get lol

OpenStudy (e.mccormick):

So now, if we leave the \(f(-8)\) alone, and remember that the \((-8)\) replaces every x, negative or not, and that \(-4\) stays negative because it does not have an x with it to cause any changes, we gt this far: \(f(-8)=-(-8)-4\implies \) \(f(-8)=-1\times (-8)-4\implies \) \(f(-8)=8-4\) What is the last step?

OpenStudy (anonymous):

I think i subtract and get 4 lol

OpenStudy (e.mccormick):

/cheer Yes

OpenStudy (anonymous):

im really smart i am lol

OpenStudy (e.mccormick):

Hey! Don't knock yourself. Like I said, this is all new. You are learning a language. It is the language of math. Before you know the rules, how can you use them to say the right things mathematically?

OpenStudy (anonymous):

so my answer would be 4? ive always struggled in math. made good grades in math but it was always hard for me to understand. i need to read more on it lol

OpenStudy (e.mccormick):

Yes, the answer is 4. And don't knock yourself. Like I said, this is all new. You are learning a language. It is the language of math. Before you know the rules, how can you use them to say the right things mathematically? Like \(\implies\). Did you know that is a word in math? It means "implies" and \(\therefore\) means therefore. Until you knew what \(+\), \(-\), \(\times\), and \(\div\) meant, you could not use them properly!

OpenStudy (anonymous):

yea seems like the more higher in math you go the higher the language goes too lol thanks for your help!

OpenStudy (anonymous):

im just checking my work to see if i have them all correct.

OpenStudy (e.mccormick):

So the new term you are learning is "function" and it comes with a whole new set of rules to follow. There are restrictions, like the vertical line test, and eventually you will learn what a one-to-one function is. Along the way, you will have to use what you already know and you will have to fit it all in together.

OpenStudy (e.mccormick):

This keeps changing too. Knwo how \(\cdot\) is used to mean \(\times\)? I am in lineear al;gebra, and at that point \(\cdot\) means dot product and \(\times\) means cross product and they are different! So yah, as you keep going it changes. Speaking of linear, I have to go to class! Hehe. Have fun here!

OpenStudy (anonymous):

lol have a great time!